OpenStudy (anonymous):

simplify tan^2 45 degrees

OpenStudy (anonymous):

sqrt2/2

OpenStudy (anonymous):



OpenStudy (anonymous):

tan^2 45 =(tan 45)^2= (1)^2= 1

OpenStudy (anonymous):

yeah i said 1